Spheres A and B have the same initial positive charge qO.The magnitude of the repulsive electric force between the two spheres is 14.4 N when separated by a distance of 0.03 m.



@



a) Find the initial q0



b)The spheres are then allowed to touch each other and then separated.Find the electric force between them if the radius of sphere A is three times that of sphere B.

1
Expert's answer
2022-01-26T17:47:18-0500

(A)F=14.4N

d=0.03m

Force

"F=\frac{kq_1q_2}{r^2}"

"14.4=\frac{9\times10^9\times q_0^2}{0.03^2}"

"1.44\times10^{-12}=q_0^2"

"q_0=1.2\times10^{-6}C"

Charge

"q_0=q_1=q_2=1.2\times10^{-6}C"

Part(b)

According to question


"\frac{q}{r}=\frac{q_0-q}{3r}"

"3q=q_0-q"

"4q=q_0"

"q=\frac{q_0}{4}"

Now force

"F=\frac{kq_1q_2}{r^2}"

"q_1=\frac{3q_0}{4},q_2=\frac{q_0}{4}"

"F=\frac{k\times\frac{3q_0}{4}\times\frac{q_0}{4}}{0.04^2}"


"F=\frac{9\times10^9\times3\times1.2\times10^{-6}\times1.2\times10^{-6}}{16\times0.0016}=2.7N"